OpenStudy (anonymous):

One of the factors of 6x^3 - 864x is 4 x2 x + 12 x - 8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I need teaching xD

OpenStudy (anonymous):

so how can I find a factor and what is a factor

ganeshie8 (ganeshie8):

first see that, it has two terms. can u pull out something common ?

OpenStudy (anonymous):

They both have x but one is x^3

ganeshie8 (ganeshie8):

yes, they have x in common also they have 6 in common

ganeshie8 (ganeshie8):

pull out the common factor 6x :- 6x^3 - 864x 6x (x^2 - 144)

ganeshie8 (ganeshie8):

fine ?

OpenStudy (anonymous):

o_O I didn't figure 6 would be a common but alright works for me :D

OpenStudy (anonymous):

lol how' 864 go to 144?

ganeshie8 (ganeshie8):

wats 6*144 ? :)

OpenStudy (anonymous):

oh I see... its 864

OpenStudy (anonymous):

ok so now what

ganeshie8 (ganeshie8):

6x^3 - 864x 6x (x^2 - 144)

ganeshie8 (ganeshie8):

is 144 a perfect square ?

ganeshie8 (ganeshie8):

12*12 = 144, so yes its a perfect square :- 6x^3 - 864x 6x (x^2 - 144) 6x(x^2-12^2)

OpenStudy (anonymous):

yeah I think

ganeshie8 (ganeshie8):

next use the identity :- a^2-b^2 = (a+b)(a-b)

ganeshie8 (ganeshie8):

wat do u get ?

OpenStudy (anonymous):

I'm not sure how to input 6x(x^2-12^2) into that equation

ganeshie8 (ganeshie8):

\(a^2-b^2 = (a+b)(a-b)\) so, \(x^2-12^2 = ?\)

OpenStudy (anonymous):

-12x^2?

ganeshie8 (ganeshie8):

let me show u once, then u wil see how to to :- \(a^2-b^2 = (a+b)(a-b)\) so, \(x^2 - 12^2 = (x+12)(x-12)\)

ganeshie8 (ganeshie8):

now, get back to oiur original problem :- 6x^3 - 864x 6x (x^2 - 144) 6x(x^2-12^2) 6x(x+12)(x-12)

ganeshie8 (ganeshie8):

see if it makes some sense

OpenStudy (anonymous):

so it would be 6x^3 - 12^3 = (x+12)(x-12)?

ganeshie8 (ganeshie8):

wat happened to the 6x before it ?

ganeshie8 (ganeshie8):

6x^3 - 12^3 = 6x(x+12)(x-12)

ganeshie8 (ganeshie8):

we say : 6, x, x+12 and x-12 are the factors

ganeshie8 (ganeshie8):

look at ur options and see they have any of these

OpenStudy (anonymous):

man this factor stuff is confusing lol

OpenStudy (anonymous):

x + 12 is an option

OpenStudy (anonymous):

Alright I think I got it :)

ganeshie8 (ganeshie8):

we need to remember few identities like the one we have used in this :- \(a^2-b^2 = (a+b)(a-b)\)

ganeshie8 (ganeshie8):

oly that identity i guess, if u are able to identify that wid ease, then its all easy :)

OpenStudy (anonymous):

yeah, can you lend me a hand on a couple more? I want to make sure I get the identity formula down :)

ganeshie8 (ganeshie8):

okay sure :)

OpenStudy (anonymous):

thanks